OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3

sam (.sam.):

4x + 2y = 12 x – y = 3 ------------------------- x=3+y 4(3+y) + 2y = 12 12+4y+2y=12 12+6y=12 6y=0 y=0 ------------------- x=3